# Exponentially weighted estimands and the exponential family: filtering, prediction and smoothing

### v1

@@ 4.3 Gaussian case: comparison to Kalman filter
The Kalman filter implies $a_{t|t-1}={\mathrm{E}}[\mu_{t}|Y_{1:t-1}]={\mathrm{E}}[\mu_{t-1}|Y_{1:t-1}]$ is given by (e.g. Harvey(89) and DurbinKoopman(12))
@@
where $K_{t}$ is called the Kalman gain and $\sigma^{2}_{\varepsilon}P_{t}={\mathrm{V}}(\mu_{t}|Y_{1:t-1})$. If $q>0$, in steady state, $K_{t}$ and $P_{t}$ converge to
$K=\frac{P}{P+1},\quad\text{and}\quad P=K+q.$
Then $K^{2}+qK-q=0$, implying $K=\frac{-q+\sqrt{q^{2}+4q}}{2}.$ In steady state, the one-step predictor is
@@
We note that if $\lambda>0$, then $n_{\lambda,t}=\frac{1-\lambda^{t}}{1-\lambda}.$ Hence it is interesting to plot
$K_{t}\times n_{\lambda,t}$
against $t$, when $P_{1|0}=\infty$, the so called diffuse initial conditions for a Kalman filter (e.g. Chapter 1 of DurbinKoopman(12)). If this product is less than one, then our recursion gives more weight on $Y_{t}$ than the Kalman filter, and subsequently, less on the past.
@@
The easiest way to see differences is to think of $q$ as small, then, for example,
$\lambda\approx 1-q^{1/2},\quad n_{\lambda,2}\approx 2-q^{1/2},\quad K_{2}\approx\frac{1}{2}-q,\quad K_{2}n_{\lambda,2}\approx 1-q^{1/2}/2.$
Figure 4 plots the product $K_{t}\times n_{\lambda,t}$ for $q\in\{0.001,0.1,0.3,1,2,10\}$ against $t$. The worst case is when $q$ is tiny, as it takes quite a large $t$ for the effect to disappear (as we move to the steady state).
@@
5 A quasi-likelihood approach to hyperparameter estimation
@@ 5.1 A quasi-likelihood interpretation
(a) $\mathbb{E}[|h(Y_{t})|]<\infty$ and denote
$\mathbb{E}[h_{t}(Y_{t})|Y_{1:t-1}]:=\mu_{t|t-1},\quad t=1,...,T.$
(b) ${\mathrm{V}}[h_{t}(Y_{t})]$ exists, the
$\Sigma_{t|t-1}:={\mathrm{V}}[h_{t}(Y_{t})|Y_{1:t-1}],\quad t=1,...,T$
and there exists a positive definite matrix $C$ such that $\Sigma_{t|t-1}-C$ is positive semidefinite for every $t$.
@@
by convexity of $\psi$ from Definition 3 and then using p.69 of BoydVandenberghe(04). The equality is only obtained iff
$\widetilde{\theta}_{t|t-1}=\theta_{t|t-1},$
that is only if the working model has the correct predictive mean. Thus under Assumptions 4(a) and 5(a) the log-likelihood process $\{l_{t}\}_{t=1}^{T}$ can be view as a quasi-likelihood process.
@@
$\langle l\rangle_{t}$ $:=\sum_{j=1}^{t}{\mathrm{V}}(l_{j}|Y_{1:j-1}),\quad t=1,...,T$
$=\sum_{j=1}^{t}\widetilde{\theta}_{j|j-1}^{\tt T}\Sigma_{j|j-1}\widetilde{\theta}_{j|j-1},$
$=\langle M-C\rangle_{t}$
@@
(a) This quasi-likelihood interpretation is not surprising, it falls in line with the storied history of, for example, generalized linear models, e.g. Wedderburn(74) and McCullaghNelder(89). More broadly quasi-likelihoods have played a vast role in modern statistics with the estimation theory going back at least to Cox(61), Huber(67), Gallant(87) and White(82).
@@ 5.2 Maximum likelihood estimation
Now turn to the maximum likelihood estimator of the hyperparameters based on the above exponential family quasi-likelihood, where the predictor comes from exponential weighting and uses a canonical exponential family frame.
@@
For a strictly stationary stochastic process $\{Y_{t}\}_{t\geq 1}$, assume that $\mathbb{E}[h(Y_{1})]$ exists and lies in $\mathcal{H}:=h(\mathcal{Y})\subseteq\mathbb{R}^{k}$. For the exponentially weighted estimands under the stable frame minimal canonical exponential family ${\tt CEF}(\theta,h,\psi)$, in steady state,
@@
Throughout we will posit that $\widetilde{\theta}_{t|t-1}(\omega)$ is infinitely differentiable for all $\omega\in\Omega^{*}$ and we write
@@
(a) Let $s_{t}:=s_{t}(\omega^{*};Y_{1:t})$, and $S_{t}=\sum_{j=1}^{t}s_{j}$. Then $\{S_{t}\}_{t=1}^{T}$ is a martingale sequence with respect to the data’s natural filtration so long as $\mathbb{E}[|s_{t}|]<\infty.$ Under the Brown(71) martingale central limit theorem the
@@
$\{\langle S,S\rangle_{T}^{-1/2}\bar{H}_{T}\}(\hat{\omega}_{\tt MLE}-\omega^{*})$ $=\langle S,S\rangle_{T}^{-1/2}S_{T}$
$\xrightarrow{D}N(0,I).$
If $\bar{H}_{T}$ is invertible, then
$W_{T}=\bar{H}^{-1}_{T}\langle S,S\rangle_{T}\bar{H}^{-1}_{T},$
is an infeasible approximation to the variance-covariance matrix of the estimator.
(c) In practice we use the estimator
$\hat{W}_{T}=H^{-1}_{T}(\hat{\omega}_{\tt MLE})[\hat{S},\hat{S}]_{T}H^{-1}_{T}(\hat{\omega}_{\tt MLE}),\quad\text{where}\quad[\hat{S},\hat{S}]_{T}=\sum_{t=1}^{T}\hat{s}_{t}\hat{s}_{t}^{\tt T},\quad\text{with}\quad s_{t}=s_{t}(\hat{\omega}_{\tt MLE};Y_{1:t}),$
as a feasible approximation to the variance-covariance matrix of the estimator. Hence this is a sandwich matrix of the tradition we see in quasi-likelihood estimation, going back to at least Cox(61) and Huber(67).
@@ 5.3 Two step alternative to the MLE
An alternative to the MLE is a two step procedure (e.g. NeweyMcFadden(94), EngleMezrich(96), FrancqHorvathZakoian(13)), which has the following structure.
@@
For a strictly stationary stochastic process $\{Y_{t}\}_{t\geq 1}$, assume $\mathbb{E}[h(Y_{1})]\in\mathcal{H}$ exists. For the exponentially weighted estimands under the stable frame minimal canonical exponential family ${\tt CEF}(\theta,h,\psi)$, let $\phi$ denote any additional static parameters from the model density. Write
@@
The two step procedure can be viewed as a method of moments estimator (Pearson(1894)), based around a random function
@@
$V_{T}^{-1/2}\bar{H}_{T}(\hat{\omega}_{\tt 2Step}-\omega^{*})$ $=V_{T}^{-1/2}S_{T}$
$\xrightarrow{D}N(0,I).$
So the infeasible variance matrix for the two step estimator is
$W_{T}=\bar{H}^{-1}_{T}V_{T}\bar{H}^{-1}_{T}.$
(c) In practice we use the estimator of the covariance matrix
$\hat{W}_{T}=H^{-1}_{T}(\hat{\omega}_{\tt 2Step})\hat{V}_{T}H^{-1}_{T}(\hat{\omega}_{\tt 2Step})^{\tt T},$
where $\hat{V}_{T}$ approximates ${\mathrm{V}}(\hat{S}_{T})$. The latter can be estimated using $T$ times a long-run variance of the time series $\hat{s}_{1},...,\hat{s}_{T}$, where $\hat{s}_{t}=s_{t}(\hat{\omega}_{\tt 2Step};Y_{1:t})$.
@@ 6 Empirical example
6.1 Dirichlet based frame: household financial situation expectations
@@
For each month, only the sample proportions (percentages) falling into each category are reported, not the underlying individual responses or sample sizes. [^4] Only for the period April 2024 to April 2025 the monthly sample sizes are given. We leave them out of our analysis.
We model the observed proportions $y_{t}=(y_{1,t},\ldots,y_{7,t})^{\tt T}$ as realizations from a time-varying seven category Dirichlet distribution
$Y_{t}\mid Y_{1:t-1}\sim\text{Dirichlet}(\widetilde{\theta}_{t|t-1}),\quad\widetilde{\theta}_{t|t-1}=(\widetilde{\theta}_{1,t|t-1},\ldots,\widetilde{\theta}_{7,t|t-1})^{\tt T},$
with concentration parameters $\widetilde{\theta}_{j,t|t-1}>0$ for all $j\in\{1,\ldots,7\}$ and all $t$. Each observation is treated as a single draw of a proportion vector and the predictor is based on a Dirichlet frame.
@@
Following the two-step quasi-likelihood procedure from Definition 5, we first estimate the centering parameter $\mathbb{E}[h(Y_{1})]$ by the sample mean
$\widehat{\mathbb{E}[h(Y_{1})]}=T^{-1}\sum_{t=1}^{T}(\log y_{1,t},\ldots,\log y_{7,t})^{\tt T}.$
Next, we fix $\widehat{\mathbb{E}[h(Y_{1})]}$ and estimate $(\alpha,\lambda)$ by maximizing the quasi-log-likelihood
$\sum_{t=1}^{T}\log f(y_{t};\widetilde{\theta}_{t|t-1}(\alpha,\lambda)).$
@@
Figure 5 presents the estimation results. Each panel shows the evolution of one expectation category over the 47-year period. The gray circles represent observed monthly proportions $y_{k,t}$. The light colored lines show the exponentially weighted predictor, the dark colored lines the exponentially weighted smoother. Both successfully track major shifts in household financial expectations. Patterns that stand out are the sharp declines in both “Continuous Increase” and “Intermittent Increase” during the 2008 financial crisis, and, more strikingly, since around 2018. This deterioration in optimism about household finances is mirrored by corresponding increases in negative expectations: “Intermittent Decline” and “Continuous Decline” roughly track their inverse patterns. The categories “Remain Unchanged” and “Don’t Know/No Answer” stay relatively stable at 15-20% and 3%, respectively, throughout the sample.
@@
Figure 5: Household financial situation expectations from the University of Michigan Survey of Consumers (January 1978 to September 2025, $T=573$ monthly observations). Categories combine responses about current versus past financial situation, and expectations for the year ahead (Cont. = Continuous, Int. = Intermittent, DK = Don’t Know, NA = No Answer): Continuous Increase means better off now and expecting better, Intermittent Increase indicates improvement in one period only, and analogously for declines. Each panel shows, for one response category, the observed monthly proportions $y_{k,t}$ (gray circles), the predictor $y_{k,t|t-1}=\widetilde{\theta}_{k,t|t-1}/\sum_{j=1}^{7}\widetilde{\theta}_{j,t|t-1}$ (light colored line), and the smoother $y_{k,t|T}=\widetilde{\theta}_{k,t|T}/\sum_{j=1}^{7}\widetilde{\theta}_{j,t|T}$ (dark colored line). Estimates: $(\widehat{\alpha},\widehat{\lambda})=(0.95,0.65).$
@@ 6.2 Monte Carlo assessment of estimator precision
To investigate the finite-sample precision of the quasi-likelihood estimator from Definition 5, we conduct a Monte Carlo study using the estimated model from Section 6.1 as the data-generating process (DGP). We generate time-varying proportion vectors following Assumption 3, that is,
$Y_{t}\mid Y_{1:t-1}\sim{\tt Dirichlet}(\theta_{t|t-1}),\quad\theta_{t|t-1}=(\theta_{1,t|t-1},\ldots,\theta_{7,t|t-1})^{\top},$
where $\theta_{t|t-1}$ evolves according to the exponentially weighted predictor (Theorem 1), denoting $(\pi^{*},\alpha^{*},\lambda^{*})$ for the hyperparameters in the DGP. We initialize each prediction path with $\widetilde{\theta}_{1|0}=(\psi^{\prime})^{-1}(\pi^{*})$ where $\pi^{*}=\widehat{\mathbb{E}[h(y_{1})]}$ is the estimated centering parameter from Section 6.1.
@@
Figure 6: Monte Carlo assessment of quasi-likelihood estimator precision for household financial expectations across varying anchoring levels and sample sizes. Rows correspond to four DGP scenarios with anchoring hyperparameters: (i) $\alpha^{*}=0.01$, (ii) $\alpha^{*}=0.30$, (iii) $\alpha^{*}=0.60$, and (iv) $\alpha^{*}=0.95$ (matching estimated household data). All scenarios use the estimated centering parameter $\pi^{*}=\widehat{\mathbb{E}[h(y_{1})]}=(-1.76,-1.41,-1.78,-1.77,-2.73,-2.23,-3.53)^{\prime}$ and discount parameter $\lambda^{*}=0.65$ from Section 6.1. Each panel displays the quasi log-likelihood surface $\ell(\alpha,\lambda)$ with the maximum quasi-likelihood estimate (red star) and DGP value (blue circle). The red contour traces the 99% confidence region. Columns show increasing sample sizes $T\in\{100,1000,5000,10000\}$.

### v2

@@ 4.3 Gaussian case: comparison to Kalman filter
The Kalman filter implies $a_{t|t-1}={\mathrm{E}}[\mu_{t}|Y_{1:t-1}]={\mathrm{E}}[\mu_{t-1}|Y_{1:t-1}]$ is given by (e.g. Harvey (1989) and Durbin and Koopman (2012))
@@
where $K_{t}$ is called the Kalman gain and $\sigma^{2}_{\varepsilon}P_{t}={\mathrm{V}}(\mu_{t}|Y_{1:t-1})$. If $q>0$, in steady state, $K_{t}$ and $P_{t}$ converge to $K=\frac{P}{P+1}$ and $P=K+q.$ Then $K^{2}+qK-q=0$, implying $K=\frac{-q+\sqrt{q^{2}+4q}}{2}.$ In steady state, the one-step predictor is
@@
We note that if $\lambda>0$, then $n_{\lambda,t}=\frac{1-\lambda^{t}}{1-\lambda}.$ Hence it is interesting to plot $K_{t}\times n_{\lambda,t}$ against $t$, when $P_{1|0}=\infty$, the so called diffuse initial conditions for a Kalman filter (e.g. Chapter 1 of Durbin and Koopman (2012)). If this product is less than one, then our recursion gives more weight on $Y_{t}$ than the Kalman filter, and subsequently, less on the past.
@@
The easiest way to see differences is to think of $q$ as small, then, for example, $\lambda\approx 1-q^{1/2}$, $n_{\lambda,2}\approx 2-q^{1/2}$, $K_{2}\approx\frac{1}{2}-q,$ and $K_{2}n_{\lambda,2}\approx 1-q^{1/2}/2.$ Figure 4 plots the product $K_{t}\times n_{\lambda,t}$ for $q\in\{0.001,0.1,0.3,1,2,10\}$ against $t$. The worst case is when $q$ is tiny, as it takes quite a large $t$ for the effect to disappear (as we move to the steady state).
@@
5 Hyperparameter estimation
@@ 5.1 A quasi-likelihood interpretation
(a) $\mathbb{E}[|h(Y_{t})|]<\infty$ and denote $\mathbb{E}[h_{t}(Y_{t})|Y_{1:t-1}]:=\mu_{t|t-1},$ where t=1,,T.t=1,...,T.
(b) ${\mathrm{V}}[h_{t}(Y_{t})]$ exists, the $\Sigma_{t|t-1}:={\mathrm{V}}[h_{t}(Y_{t})|Y_{1:t-1}],$ for t=1,,Tt=1,...,T and there exists a positive definite matrix $C$ such that $\Sigma_{t|t-1}-C$ is positive semidefinite for every $t$.
@@
by convexity of $\psi$ from Definition 3 and then using p.69 of Boyd and Vandenberghe (2004). The equality is only obtained iff $\widetilde{\theta}_{t|t-1}=\theta_{t|t-1},$ that is only if the working model has the correct predictive mean. Thus under Assumptions 4(a) and 5(a) the log-likelihood process $\{l_{t}\}_{t=1}^{T}$ can be view as a quasi-likelihood process.
@@
$\langle l\rangle_{t}$ $:=\sum_{j=1}^{t}{\mathrm{V}}(l_{j}|Y_{1:j-1})=\sum_{j=1}^{t}\widetilde{\theta}_{j|j-1}^{\tt T}\Sigma_{j|j-1}\widetilde{\theta}_{j|j-1}=\langle M-C\rangle_{t},\quad t=1,...,T$
@@
(a) This quasi-likelihood interpretation is not surprising, it falls in line with the storied history of, for example, generalized linear models, e.g. Wedderburn (1974) and McCullagh and Nelder (1989). More broadly quasi-likelihoods have played a vast role in modern statistics with the estimation theory going back at least to Cox (1961), Huber (1967), Gallant (1987) and White (1982).
@@ 5.2 Maximum likelihood estimation
Now turn to the maximum likelihood estimator of the hyperparameters based on the above exponential family quasi-likelihood, where the predictor comes from exponential weighting and uses a CEF frame.
@@
For a strictly stationary stochastic process $\{Y_{t}\}_{t\geq 1}$, assume that $\mathbb{E}[h(Y_{1})]$ exists and lies in $\mathcal{H}:=h(\mathcal{Y})\subseteq\mathbb{R}^{k}$. For the exponentially weighted estimands under the stable frame minimal ${\tt CEF}(\theta,h,\psi)$, in steady state,
@@
Throughout we will posit that $\widetilde{\theta}_{t|t-1}(\omega)$ is infinitely differentiable for all $\omega\in\Omega^{*}$ and write
@@
(a) Let $s_{t}:=s_{t}(\omega^{*};Y_{1:t})$, and $S_{t}=\sum_{j=1}^{t}s_{j}$. Then $\{S_{t}\}_{t=1}^{T}$ is a martingale sequence with respect to the data’s natural filtration so long as $\mathbb{E}[|s_{t}|]<\infty.$ Under the M (1971) martingale central limit theorem the
@@
$\{\langle S,S\rangle_{T}^{-1/2}\bar{H}_{T}\}(\hat{\omega}_{\tt MLE}-\omega^{*})$ $=\langle S,S\rangle_{T}^{-1/2}S_{T}\xrightarrow{D}N(0,I).$
If $\bar{H}_{T}$ is invertible, then $W_{T}=\bar{H}^{-1}_{T}\langle S,S\rangle_{T}\bar{H}^{-1}_{T},$ is an infeasible approximation to the variance-covariance matrix of the estimator.
(c) In practice we use the estimator $\hat{W}_{T}=H^{-1}_{T}(\hat{\omega}_{\tt MLE})[\hat{S},\hat{S}]_{T}H^{-1}_{T}(\hat{\omega}_{\tt MLE}),$ where $\hat{S},\hat{S}]_{T}=\sum_{t=1}^{T}\hat{s}_{t}\hat{s}_{t}^{\tt T},$ with $s_{t}=s_{t}(\hat{\omega}_{\tt MLE};Y_{1:t}),$ as a feasible approximation to the variance-covariance matrix of the estimator. Hence this is a sandwich matrix of the tradition we see in quasi-likelihood estimation, going back to at least Cox (1961) and Huber (1967).
@@ 5.3 Two step alternative to the MLE
An alternative to the MLE is a two step procedure (e.g. Newey and McFadden (1994), Engle and Mezrich (1996), Francq et al. (2013)), which has the following structure.
@@
For a strictly stationary stochastic process $\{Y_{t}\}_{t\geq 1}$, assume $\mathbb{E}[h(Y_{1})]\in\mathcal{H}$ exists. For the exponentially weighted estimands under the stable frame minimal ${\tt CEF}(\theta,h,\psi)$, let $\phi$ denote any additional static parameters from the model density. Write
@@
The two step procedure can be viewed as a method of moments estimator (Pearson, 1894), based around a random function
@@
$V_{T}^{-1/2}\bar{H}_{T}(\hat{\omega}_{\tt 2Step}-\omega^{*})$ $=V_{T}^{-1/2}S_{T}\xrightarrow{D}N(0,I).$
So the infeasible variance matrix for the two step estimator is $W_{T}=\bar{H}^{-1}_{T}V_{T}\bar{H}^{-1}_{T}.$
(c) In practice we use the estimator of the covariance matrix $\hat{W}_{T}=H^{-1}_{T}(\hat{\omega}_{\tt 2Step})\hat{V}_{T}H^{-1}_{T}(\hat{\omega}_{\tt 2Step})^{\tt T},$ where $\hat{V}_{T}$ approximates ${\mathrm{V}}(\hat{S}_{T})$. The latter can be estimated using $T$ times a long-run variance of the time series $\hat{s}_{1},...,\hat{s}_{T}$, where $\hat{s}_{t}=s_{t}(\hat{\omega}_{\tt 2Step};Y_{1:t})$.
@@ 6 Empirical example
6.1 Dirichlet based frame: household financial situation
@@
For each month, only the sample proportions (percentages) falling into each category are reported, not the underlying individual responses or sample sizes. [^3] While monthly sample sizes are available, our framework would require within-month i.i.d. responses, which is unlikely given the clustered sampling design. We leave them out of our analysis.
We model the observed proportions $y_{t}=(y_{1,t},\ldots,y_{7,t})^{\tt T}$ as realizations from a time-varying seven category Dirichlet distribution $Y_{t}\mid Y_{1:t-1}\sim\text{Dirichlet}(\widetilde{\theta}_{t|t-1}),$ with $\widetilde{\theta}_{t|t-1}=(\widetilde{\theta}_{1,t|t-1},\ldots,\widetilde{\theta}_{7,t|t-1})^{\tt T},$ with concentration parameters $\widetilde{\theta}_{j,t|t-1}>0$ for all $j\in\{1,\ldots,7\}$ and all $t$. Each observation is treated as a single draw of a proportion vector. The predictor is based on a Dirichlet frame.
@@
Following the two-step quasi-likelihood procedure from Definition 5, we estimate the centering parameter $\mathbb{E}[h(Y_{1})]$ by the sample mean $\widehat{\mathbb{E}[h(Y_{1})]}=T^{-1}\sum_{t=1}^{T}(\log y_{1,t},\ldots,\log y_{7,t})^{\tt T}.$ Next, we fix $\widehat{\mathbb{E}[h(Y_{1})]}$ and estimate $(\alpha,\lambda)$ as maximizing $\sum_{t=1}^{T}\log f(y_{t};\widetilde{\theta}_{t|t-1}(\alpha,\lambda)).$
@@
Figure 5 presents the estimation results. Each panel shows the evolution of one expectation category over the 47-year period. The gray circles represent observed monthly proportions $y_{k,t}$. The light colored lines show the exponentially weighted predictor, the dark colored lines the exponentially weighted smoother. Both track major shifts in household financial expectations. There are the sharp declines in both “Continuous Increase” and “Intermittent Increase” during the 2008 financial crisis, and, more strikingly, since around 2018. This deterioration in optimism about household finances is mirrored by corresponding increases in negative expectations: “Intermittent Decline” and “Continuous Decline” roughly track their inverse patterns. The categories “Remain Unchanged” and “Don’t Know/No Answer” stay relatively stable at 15-20% and 3%, respectively, throughout the sample.
@@
Figure 5: Household financial situation: Uni. Michigan Survey of Consumers (Jan 1978 to Sept 2025). Categories combine current versus past financial situation, & expectations for the year ahead (Cont. = Continuous, Int. = Intermittent, DK = Don’t Know, NA = No Answer): Continuous Increase means better off now & expecting better, Intermittent Increase indicates improvement in one period only. Each panel shows, for one response category, the observed monthly proportions $y_{k,t}$ (gray circles), the predictor $y_{k,t|t-1}=\widetilde{\theta}_{k,t|t-1}/\sum_{j=1}^{7}\widetilde{\theta}_{j,t|t-1}$ (light line) & the smoother $y_{k,t|T}=\widetilde{\theta}_{k,t|T}/\sum_{j=1}^{7}\widetilde{\theta}_{j,t|T}$ (dark line). Estimates: $(\widehat{\alpha},\widehat{\lambda})=(0.95,0.65).$
@@ 6.2 Monte Carlo assessment of estimator precision
To investigate the finite-sample precision of the quasi-likelihood estimator from Definition 5, we conduct a Monte Carlo study using the estimated model from Section 6.1 as the data-generating process (DGP). We generate time-varying proportion vectors following Assumption 3, that is, $Y_{t}\mid Y_{1:t-1}\sim{\tt Dirichlet}(\theta_{t|t-1}),$ where $\theta_{t|t-1}=(\theta_{1,t|t-1},\ldots,\theta_{7,t|t-1})^{\top},$ where $\theta_{t|t-1}$ evolves according to the exponentially weighted predictor (Theorem 1), denoting $(\pi^{*},\alpha^{*},\lambda^{*})$ for the hyperparameters in the DGP. We initialize each prediction path with $\widetilde{\theta}_{1|0}=(\psi^{\prime})^{-1}(\pi^{*})$ where $\pi^{*}=\widehat{\mathbb{E}[h(y_{1})]}$ is the estimated centering parameter from Section 6.1.
@@
Figure 6: Monte Carlo: quasi-likelihood estimator precision for household financial expectations. Rows: anchoring hyperparameters: (i) $\alpha^{*}=0.01$, (ii) $\alpha^{*}=0.30$, (iii) $\alpha^{*}=0.60$, and (iv) $\alpha^{*}=0.95$ (matching estimated household data). All scenarios use the estimated centering parameter $\pi^{*}=\widehat{\mathbb{E}[h(y_{1})]}=(-1.76,-1.41,-1.78,-1.77,-2.73,-2.23,-3.53)^{\prime}$ and discount parameter $\lambda^{*}=0.65$ from Section 6.1. Display: quasi log-likelihood surface $\ell(\alpha,\lambda)$ with the maximum quasi-likelihood estimate (red star) and DGP value (blue circle). The red contour traces the 99% confidence region. Columns: $T\in\{100,1000,5000,10000\}$.